ARGOS Commissioning, April 22 to May 3, 2015

Players: ARGOS team

Report from Sebastian

  • In a sad combination of a bridged interlock and communication issues, the LBC camera has been deployed into our calibration arm, that massively bended. Luckily the optics pieces where not damaged. In a huge effort Wolfgang and crew exchanged the broken gusset with the SX side of the telescope and aligned the unit within three days, an action that usually takes a week.

  • One of the controllers in the NGS AO has been broken already before we started and left us without the tech viewer and without modulation on the pyramid sensor. With the extreme tiny field of the NGS APD sensor, and the pyramid, the finding of the star went into a time consuming puzzle. Finally this prevented us as well to test the APD sensors on the very dim stars. Nevertheless we experienced that modulation is not a must for the pyramid.

  • In the first night the adaptive secondary mirror decided to stop operation, riping every few minutes without reason. The investigated extended more or less over all night. Supplying better cooling to one of the Adsec electronic racks, and letting the system rest over day made then the problem disappear, but it is still unclear if this was just an accidently coincidence.

  • The software has grown quite mature, bringing us close to an automatic acquisition and AO loop closure. We have been able to send presets and offsets from LUCI, seeing them received by the telescope and the wavefront sensor. Within a next step, that we finally could not test anymore, this will enable observing scripts to be executed.

  • The NGS APD tip tilt system is tracking nicely and stable, and it looks like that this works as well with very low counts on the unit. A quantification needed to be postponed, due to the broken BCU in the FLAO, since a co-pointing between APD and Pyramid was not possible.

  • The vibration compensation system seems to do the right thing, but just checked by eye til now. A measurement on the WFS was weathered out yesterday.

  • Finally we have seen the first three nights with excellent weather conditions and good seeing. Apart from the lost first one, we therefore could do some performance tests and learn quite a lot about the system.

  • While most of the time spend with very technical tests, one of the GLAO test fields happened to be filled with NGC6384. This an image showing a J H composite with some contrast scaling applied: NGC6384_jh. The total integration time was 16min for each band. In the individual images we find a FWHM in H between 0.22 arcsec close to the tilt star, and up to 0.33 arcsec in the edges of the field, and 0.34 to 0.4 arcsec in J-band. As said the seeing conditions where pretty good during that night, with the DIMM measuring around 0.85, and the open loop performance around 0.45 in H. There is for sure quite a long way to fully understand all details, but nevertheless this is showing that we are not too far from our predictions.

  • In total there is still a huge amount of work to be done for hard and especially software to arrive at the final system, so we keep on working and are looking forward for the next steps.

TBAD operation in parallel to Spotters

  • This was the first time TBAD was used on sky during a laser run, connected to the ARGOS Interlock System.

  • TBAD detections:
    • Night 1 (April 30 UTC): 3 closures called by spotters. 1 detected by TBAD (6:34 UTC, AA1441, AC2A33)
    • Night 2 (May 1 UTC): 2 closures called by spotters. 1 detected by TBAD (6:10 UTC, AA1042, AC90A8)
    • Night 3 (May 2 UTC): No closure called by spotters. 1 detected by TBAD (8:06 UTC, AA154, ACF975)
    • Night 4 (May 3 UTC): No closures (less than 1 hour of laser propagation, rest of time closed due to closed)
    • Night 5 (May 4 UTC): Telescope closed due to weather.

Time (UTC) Spotter TBAD ID AZ EL Notes
2015-04-30
02:24 - X AD0822 83.0 52.3 Not lasing. Spotters not out yet.
02:38 - X ACA788 84.4 55.0 Not lasing. Spotters not out yet.
02:47 - X A5F5BF 85.5 57.1 Not lasing. Spotters not out yet.
06:08 X - AD3997 275.3 66.9 No TBAD closure.
06:34 X X AC2433 277.5 61.4  
06:44 X - A471E0 278.4 59.3 No TBAD closure.
2015-05-01
03:11 X X A81F50 88.9 76.5 Aircraft in area at clearance request.
03:20 X - AC97E4 90.0 78.4 No TBAD closure.
06:10 X X AC90A8 275.7 65.8  
2015-05-02
02:32 - X AD247C 90.0 90.0 Not lasing. Spotters not out yet.
04:35 - X AAA4BE 33.4 57.4 Not lasing. Spotters had been sent inside for clouds.
08:05 - X ACF975 123.8 50.9 TBAD closure. Missed by spotters.

-- %USERSIG{GustavoRahmer - 2015-05-07}%

Comments

 
Topic revision: r3 - 15 Jul 2015, GustavoRahmer
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback